OpenStudy (anonymous):

Can someone show me what this looks like together on a single graph? Please help. 7x + 2y = –6 6x – 7y = 21

OpenStudy (anonymous):

https://www.desmos.com/calculator go here and plug them in but for your information does aren't in slope formula.

OpenStudy (anonymous):

Did this help?

OpenStudy (anonymous):

Is it for Systems of Equations?

OpenStudy (anonymous):

yes but you may want does in slope form y=mx + b

OpenStudy (anonymous):

it helps

OpenStudy (anonymous):

2y=-7x-6

OpenStudy (anonymous):

yes keep going

OpenStudy (anonymous):

To graph things together, there's an awesome program called GeoGebra that's free.

OpenStudy (anonymous):

-7y=-6x+21

OpenStudy (anonymous):

i think my online school teacher mentioned a program like that, just couldnt ever find it.

OpenStudy (anonymous):

y=6/7x+21 is the equation and yes geogebra is a good program

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

slope formula for the equation that your were doing is y=6/7x+21

OpenStudy (jdoe0001):

$$ y = -3 -\frac{7x}{2}\\ y = \frac{6x}{7}-3 $$

OpenStudy (anonymous):

... yes does are the equations sorry

OpenStudy (anonymous):

i dont need slope formula. i found the right answers.

OpenStudy (anonymous):

and i did another question as well.